Solve and graph 2x=5y-11 and 3x=2y
Possible solutions to equation 1 (0,2.5) and (-5.5,0)
Possible solutions to equation 2 (0,0) and (2,3)
Note if Independent, dependent, or inconsistent

Answer by ewatrrr(24785)


Hi
2x=5y-11 OR y = (2/5)x + 11/5 (0,2.2) and (-5.5,0)
and 3x=2y OR y = (3/2)x (0,0) and (2,3)
Graphing indicates one solution to this system (2,3)
(2/5)x + 11/5 = (3/2)x
22/10 = 15/10x-4/10x
22/10 = 11/10x x = 2 and y = 3 [(3/2)2]
System is Independent